Find each probability if you pick two marbles without replacing the first
4 green. 2 red. 3 yellow
P(yellow, then not yellow)

Answer by stanbon(75887)

Find each probability if you pick two marbles without replacing the first
4 green. 2 red. 3 yellow
P(yellow, then not yellow) = (3/9)(6/8) = 1/4
==============
Cheers,
Stan H.
==============